Solve by elimination method

x + 6y = 55
-x + 3y = 17

Answer by stanbon(75887)

Solve by elimination method
x + 6y = 55
-x + 3y = 17
--------------------
Add to get:
9y = 72
y = 8
------
Since x +6*8 = 55
x = 7
=======
Cheers,
Stan H.